# On Schultz’s generalization of Borweins’ cubic identity

### v1

@@ 1. Introduction
The following theorems provide generalizations of (5.1) and (5.2), respectively.
Theorem 5.1.
Suppose $b$ and $d$ are integers satisfying 2d<b<2d-2d<b<2d. Then
Theorem 5.2.
Note that (1.6) follows from (5.3) when $d=1$ and $b=0$. We derive Theorem 5.1 from the following result.
@@
Similarly, we obtain Theorem 5.2 from the following.
Theorem 5.4.
Let
and
Let $0<\alpha<4\beta$ so that both double sums converge. Then
Moreover,
Proof of Theorem 5.3.
Every integer pair $(m,n)\in\mathbf{Z}^{2}$ corresponds uniquely to a pair $(k,\ell)\in\mathbf{Z}^{2}$ satisfying $k\equiv\ell\pmod{2}$ via the transformation
@@
$\sum_{m,n\in\mathbf{Z}}F(m,n)$ $=\sum_{\begin{subarray}{c}k,\ell\in\mathbf{Z}\\ k\equiv\ell\pmod{2}\end{subarray}}F\left(\frac{k+\ell}{2},\frac{k-\ell}{2}\right)$
@@
and the proof of Theorem 5.3 is complete.
Proof of Theorem 5.1.
Set $x=e^{2i(u+v)}$ and $y=e^{2i(u-v)}$ in (5.3). With these substitutions, (5.3) becomes equivalent to (5.8). Since (5.8) holds, the theorem follows immediately. ∎
Next, we show that (5.8) is a two-variable generalization of (5.1).
From (5.14), we find that
@@
where we have used [7, p.58]
@@
Proof of Theorem 5.4.
and
Therefore, by applying (5.7),
which is (5.11). Identity (5.12) also follows easily from (5.8) by nothing that
$=\widehat{A}_{\alpha,\beta}^{2}(u,v+\pi/2|\tau)-\widehat{C}_{\alpha,\beta}^{2}(u,v+\pi/2|\tau).$
Proof of Theorem 5.2.
Set $x=e^{4iu}$ and $y=e^{2i(u+v)}$ in (5.4). With these substitutions, (5.4) becomes equivalent to (5.12). Since (5.12) holds, the theorem follows immediately. ∎
Finally, we show that (5.12) is a two variable generalization of [8, (1.6)]. Comparing (5.12) and [8, (1.6)], we see that it suffices to show
@@
Since it is shown in [8, (3.2)] that
@@
6. Cubic analogues of (1.1)
It is natural to ask if there are cubic analogues of (5.1) and (5.2). This question leads us to the following new identities of Borweins-type.
@@
We note that (6.1) is the cubic analogue of (5.1) in the case $b=d=1$ and (6.2) is the cubic analogue of (5.2) in the case $b=3$ and $d=1$. These identities are different ways of expressing (1.1) but establishing their connections with (1.1) turns out to be very interesting.
@@
$R(m+n\gamma)=m$
and
$I(m+n\gamma)=n,$
@@
$\mathcal{L}=\{m+n\gamma|m,n\in\mathbf{Z}\}\quad\text{and}\quad\mathcal{L}_{j}=\{-2n-m+j+(2m+n)\gamma|m,n\in\mathbf{Z}\},$
@@
$\sum_{m,n\in\mathbf{Z}}q^{3m^{2}+3n^{2}+3mn+3m+3n+1}=\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ m\equiv n+1\!\!\!\!\!\pmod{3}\end{subarray}}q^{m^{2}+mn+n^{2}}.$
The solutions to $m\equiv n+1\pmod{3}$ are $(m,n)\equiv(1,0),(-1,1)$ and $(0,-1)\pmod{3}$. Therefore,
$\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ m\equiv n+1\!\!\!\!\!\pmod{3}\end{subarray}}q^{m^{2}+mn+n^{2}}=C_{1,0}+C_{0,-1}+C_{-1,1},$
@@
$C_{i,j}=\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ (m,n)\equiv(i,j)\!\!\!\!\!\pmod{3}\end{subarray}}q^{m^{2}+mn+n^{2}}.$
@@
    Identity (6.3) can also be proved by writing $S$ as S=ω^-1S_-1+S_0+ωS_1, where S_j = ∑_m,n∈Zm-n≡j (mod3) q^m^2+mn+n^2+m+n. It can be shown that that $S_{-1}=S_{0}=S_{1}$ so that S=(ω^-1+1+ω) S_0 = 0 since $1+\omega+\omega^{-1}=0.$ We leave the details to the reader.
@@
    Identity (6.5) is not new. The identity and its proof can be found in [25], [13, Lemma 3.13] and [11, (1.11)].
@@
which is the cubic analog of [8, eq. (1.5)] with $b=d=1$.
@@ 7. Concluding remarks
In this article, we have shown that $A(x,y|\tau)$ and $C(x,y|\tau)$ are coefficients of certain theta series identities such as (2.12). We then use (2.12) to prove (1.4). We have also connected (1.4) to identity connected to the study of Macdonald identities. Using transformation formulas from classical texts, we derive new analogues of (1.4). Motivated by (1.4) and identities discovered in [8], we found certain $X,Y,Z,W$, which are two-variable extension of series discovered in [8], that satisfy
$X^{2}+Y^{2}=Z^{2}+W^{2},$
We end the article with several new series $A,B,C$ that satisfy
$A^{3}+B^{3}=C^{3}.$
All the identities we found in that section are different ways of expressing (1.1).

### v2

@@ 5. Generalizations of Schult’z identity
Proof.
Every integer pair $(m,n)\in\mathbf{Z}^{2}$ corresponds uniquely to a pair $(k,\ell)\in\mathbf{Z}^{2}$ satisfying $k\equiv\ell\;\allowbreak(\operatorname{mod}2)$ via the transformation
@@
$\sum_{m,n\in\mathbf{Z}}F(m,n)$ $=\sum_{\begin{subarray}{c}k,\ell\in\mathbf{Z}\\ k\equiv\ell\;\allowbreak(\operatorname{mod}2)\end{subarray}}F\left(\frac{k+\ell}{2},\frac{k-\ell}{2}\right)$
@@
and the proof of Lemma 5.1 is complete. ∎
Corollary 5.1 (Generalization of (5.1)).
Suppose $b$ and $d$ are integers satisfying 2d<b<2d-2d<b<2d. Then
Proof.
Identity (5.12) follows by letting
$x=e^{4iu}\,\,\text{and}\,\,y=e^{2i(u+v)}$
in (5.6). ∎
Note that Theorem 1.2 follows from Corollary 5.1 when $d=1$ and $b=0$.
Theorem 5.2.
Let
and
Let $0<\alpha<4\beta$ so that both double sums converge. Then
Moreover,
Proof.
and
Therefore, by applying (5.5),
which is (5.15). Identity (5.16) also follows easily from (5.6) by noting that
$=\widehat{A}_{\alpha,\beta}^{2}(u,v+\pi/2|\tau)-\widehat{C}_{\alpha,\beta}^{2}(u,v+\pi/2|\tau).\qed$
Corollary 5.2 (Generalization of (5.2)).
Proof.
Identity (5.17) follows by letting
$x=e^{4iu}\,\,\text{and}\,\,y=e^{2i(u+v)}$
in (5.16). ∎
Remark 5.1.
The identity (5.6) is a two-variable generalization of (5.1) and can be shown as follow: From (5.8), we find that
@@
where we have used [9, p.58]
@@
Remark 5.2.
Identity (5.16) is a two variable generalization of [10, (1.6)] and we present the proof as follow: Comparing (5.16) and [10, (1.6)], we see that it suffices to show
@@
Since it is shown in [10, (3.2)] that
@@
6. Cubic Borweins-type identities
It is natural to ask if there are cubic analogues of (5.1) and (5.2). This question leads us to the following new identities of Borweins type.
@@ 6. Cubic Borweins-type identities
We note that (6.1) is the cubic analogue of (5.1) in the case $b=d=1$ and (6.2) is the cubic analogue of (5.2) in the case $b=3$ and $d=1$. These identities are different ways of expressing (1.1) but establishing their connections with (1.1) turns out to be very interesting on its own.
@@
$R(m+n\gamma)=m\quad\text{and}\quad I(m+n\gamma)=n,$
@@
$\mathcal{L}=\{m+n\gamma\mid m,n\in\mathbf{Z}\}\quad\text{and}\quad\mathcal{L}_{j}=\{-2n-m+j+(2m+n)\gamma\mid m,n\in\mathbf{Z}\},$
@@
$\sum_{m,n\in\mathbf{Z}}q^{3m^{2}+3n^{2}+3mn+3m+3n+1}=\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ m\equiv n+1\;\allowbreak(\operatorname{mod}3)\end{subarray}}q^{m^{2}+mn+n^{2}}.$
The solutions to $m\equiv n+1\;\allowbreak(\operatorname{mod}3)$ are $(m,n)\equiv(1,0),(-1,1)$ and $(0,-1)\;\allowbreak(\operatorname{mod}3)$. Therefore,
$\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ m\equiv n+1\;\allowbreak(\operatorname{mod}3)\end{subarray}}q^{m^{2}+mn+n^{2}}=C_{1,0}+C_{0,-1}+C_{-1,1},$
@@
$C_{i,j}=\sum_{\begin{subarray}{c}m,n\in\mathbf{Z}\\ (m,n)\equiv(i,j)\;\allowbreak(\operatorname{mod}3)\end{subarray}}q^{m^{2}+mn+n^{2}}.$
@@
    Identity (6.3) can also be proved by writing $S$ as S=ω^-1S_-1+S_0+ωS_1, where S_j = ∑_m,n∈Zm-n≡j  (mod3) q^m^2+mn+n^2+m+n. It can be shown that that $S_{-1}=S_{0}=S_{1}$, so that S=(ω^-1+1+ω) S_0 = 0 since $1+\omega+\omega^{-1}=0$. We leave the details to the reader.
@@
    Identity (6.5) is not new. The identity and its proof can be found in [29], [16, Lemma 3.13] and [13, (1.11)].
@@
which is the cubic analog of [10, eq. (1.5)] with $b=d=1$.
@@ 7. Concluding remarks
In this article, we have shown that $A(x,y|\tau)$ and $C(x,y|\tau)$ are coefficients of certain theta series identities such as (2.10). We then use (2.10) to prove (1.4). We have also connected (1.4) to identity connected to the study of Macdonald identities. Using transformation formulas from classical texts, we derive new analogues of (1.4). The function $A(x,y|\tau)$ and its variations have appeared before the work of Schultz. For more details, see [2], [8], [6] and [15].
In Section 5, we discover infinitely many theta series $X,Y,Z,W$ associated with the binary quadratic forms $bm^{2}+bmn+dn^{2}$ and $dm^{2}+bmn+dn^{2}$ with $|b|<2d$, that fit the relation
$X^{2}+Y^{2}=Z^{2}+W^{2}.$
We are, however, unable to find theta series $X,Y,Z,W$ that satisfy
$X^{3}+Y^{3}=Z^{3}+W^{3}$ (7.1)
besides those that we have presented here. In particular, we are not able to find theta series satisfying (7.1) which are not associated with the binary quadratic form $m^{2}+mn+n^{2}.$
In Section 6, we derive new series of one-variable $A,B,C$ that satisfy
$A^{3}+B^{3}=C^{3}.$ (7.2)
Although we have infinitely many triplets of series of one-variable $A,B,C$ satisfying
$A^{2}+B^{2}=C^{2},$
(6.1) and (6.2) are the only solutions we found for (7.2) and these identities are different ways of expressing (1.1). We do not know if there are theta series of one-variable that satisfy (7.2) which are not associated with the binary quadratic form $m^{2}+mn+n^{2}$.
Acknowledgments. We thank the anonymous referees for their healthy criticism that helped to improve the original version of this manuscript. We also thank Siew Lian Tan for her help in improving the presentation of the article.